OpenStudy (anonymous):

Okay, I'm doing problems on Non-Linear Systems of Equations and COMPLETELY LOST! Help? y=x^2+10x+11 y=x^2+x-7

OpenStudy (anonymous):

x^2+10x+11=x^2+x-7 10x+11=x-7 9x=-18 x=-2

OpenStudy (anonymous):

I don't think that's right, what happen to the x^2?

OpenStudy (anonymous):

A plot generated by Mathematica 8 is attached. Observe the x coordinate where the curves cross each other.

OpenStudy (anonymous):

Okay, so the x coordinate of the answer is negative two?

OpenStudy (anonymous):

Yes.